# itemizestditemize

### v1

@@ 1 Introduction
    GPU speedup: cuADMM. At each iteration, sGS-ADMM alternates between (a) performing projection onto positive semidefinite (PSD) cones, and (b) solving sparse linear systems. Due to the existence of many small-to-medium-scale PSD cones in the SDP relaxation, we engineer a highly optimized GPU-based implementation of sGS-ADMM in CUDA, named cuADMM and achieves up to 10×10$\times$10 × speedup compared to existing ADMM solvers such as cdcs [77] and sdpnal [76]. cuADMM is the first ADMM-based SDP solver that runs in GPUs and we demonstrate its ability to solve five trajectory optimization applications (inverted pendulum, cart-pole, vehicle landing, flying robot, and car back-in with obstacle avoidance): cuADMM solves them with optimality certificates (suboptimality below $1\%$) in minutes (when other solvers take hours and run out of memory) and seconds (when others take minutes).
@@ 2.1 Polynomial Optimization with Chain-like Sparsity
then the following POP is said to admit chain-like sparsity
@@ 2.2 The Sparse Moment-SOS Hierarchy
Theorem 2.1 provides a way to check the tightness of the relaxation and extract globally optimal solutions $z^{\star}$ of the nonconvex POP (5). In fact, condition (iii) guarantees the optimal TMS $\varphi^{\star}$ admits a representing Dirac measure supported on the unique optimal solution $z^{\star}$. Therefore, one can extract $z^{\star}$ from the entries of $\varphi^{\star}$ corresponding to degree-one monomials. Note that condition (iii) is a special case of the flat extension condition [45, 55], which states that even when the rank of $M_{k}^{\star}$ is larger than one, say $\mathrm{rank}\left(M_{k}^{\star}\right)=r$, as long as $r$ is consistent with the rank of a submatrix of $M_{k}^{\star}$, then global optimality can be certified and exactly $r$ minimizers of the POP can be extracted. We do not detail this setup because in our applications $\mathrm{rank}\left(M_{k}^{\star}\right)$ is always one (up to numerical precision).
@@ 2.3 Faster Conversion to Standard SDP
Solving the moment relaxation (10) is typically done by converting (10) into a standard linear conic optimization problem in the format of sedumi [62] or mosek [6]. The standard way to convert (10) into a conic program follows [73, §5.7.1] (e.g., as in tssos, sostools, yalmip), which will generate a conic program with not only PSD variables but also free (unconstrained) variables.
We advocate for a different conversion standard [73, §5.7.2] that generates only PSD variables and the resulting semidefinite program admits the same chain-like structure as the POP (cf. Fig. 1).
@@
Linear Objective $\left\langle C,X\right\rangle$. Recall the $k$-th clique contains variables $z(I_{k})=(x_{k-1},u_{k-1},x_{k})$. Set $k=2$, with relaxation order $\kappa=2$, the moment matrix
@@ 3 Algorithmic Backbone: sGS-ADMM
sGS-ADMM can be seen as a semi-proximal ADMM method applied to the Augmented Lagrangian of (22). Without going deep into the theory, and denoting $\Pi_{\Omega_{+}}(\cdot)$ as the projection onto $\Omega_{+}$, we present sGS-ADMM in Algorithm 1. Global convergence of sGS-ADMM for the SDP pair (13)-(22) is established in [43, 14].
@@ 5 Experiments
Baselines. For interior point methods, we choose the commercial solver mosek [6] as our baseline. For first-order methods, we choose the state-of-the-art general SDP solvers cdcs [77] and sdpnal [76] (both based on ADMM). [^3] We have tested scs [56] but found its Matlab version to always generate wrong results for multiple-block SDP problems. For first-order methods, tol is set as $10^{-4}$ in Algorithm 1.
@@
Summary Results. For the pendulum problem, we tested the performance of four SDP solvers on 100 initial states generated by a dense grid. For the other four experiments, we assessed cuADMM’s performance on 10 randomly generated initial states. Due to the significantly longer solving times of the other three solvers, we evaluated their performance on only one random initial state that cuADMM could solve, with a maximum running time set at 5 hours. Table 1 presents the problem scale, suboptimality gaps, and total running times. It shows that (1) the mean and median suboptimality gaps from cuADMM are either well below $10^{-2}$ (in all cases except for the car back-in) or near $10^{-2}$ (for the car back-in); (2) mosek can solve the pendulum and cart-pole problems very well, but cannot scale to other problems; (3) Compared with other ADMM-based solvers, cuADMM achieves up to 10×10$\times$10 × speedup in large-scale SDP problems (with $m>5\times 10^{5}$); In fact, the other ADMM solvers cannot produce $<1\%$ suboptimality certificates with the time presented in Table 1; (4) With warmstart techniques, cuADMM can solve the pendulum problem to global optimality with an average cost of 0.66 seconds.
@@
mosek cdcs sdpnal cuADMM
Pendulum $4$ $1$ $55$ $30$ $10$ $47351$ $\bm{-2.58}$ ($\bm{-2.93}$) $9.5$s $110.3$s $264.4$s cold: $31.2$s ($14.6$s)
$\bm{-2.71}$ ($\bm{-2.97}$) warm: $\bm{0.66}$s ($\bm{0.62}$s)
@@
Warmstarts with kNN. To enable data-driven methods for warmstarts, we collect $60\times 120$ mosek solutions in the state space $(\theta,\dot{\theta})\in[0,\pi]\times[-5,5]$ off-line. Given a new initial state $(\theta_{0},\dot{\theta}_{0})$, we use Delaunay Triangulation [40] to search for three nearest neighbors among the mosek’s solutions. A convex combination of these three solutions is treated as the initial solution for cuADMM.
Comparison with baselines. In the state space $(\theta,\dot{\theta})\in[0,\pi]\times[-5,5]$, we sample $10\times 10$ initial states with a dense grid. Fig. 4 bottom shows the performance of different solvers. In almost all experiments, mosek achieved suboptimality gaps as low as $10^{-7}$. Among the first-order solvers, cdcs struggled to get low dual infeasibility $\eta_{d}$, resulting in significantly higher suboptimality gaps. While sdpnal shows reliable performance, its per-iteration cost is $7$ times higher than cuADMM. With warmstarts, cuADMM solves the pendulum problem in less than $0.7$s on average, achieving 10×10$\times$10 × speedup compared to other solvers. $10\%$ initial states are hard to solve for all solvers, which form a mysterious spiral line as discussed in [24]. See §0.E.1 for a detailed discussion.
@@
Figure 4: Top: MPC of inverted pendulum with kNN warmstarts. Bottom: performance of five SDP solvers in inverted pendulum.
@@ 0.E.1 Inverted Pendulum
Hyepr-parameters. Denote $x_{k}$ as $\left[r_{c,k}\;r_{s,k}\;f_{s,k}\;f_{s,k}\right]\in{{\mathbb{R}}^{4}}$. In all experiments, $m=1,l=1,b=0.1,g=9.8$ and $N=30,\Delta t=0.1,f_{c,\min}=0.5,u_{\max}=5$. $P_{f}$ in (34) is set to $1$. The number of localizing matrices is $60$. For first-order methods, maxiter is set to $10000$.
@@
Refer to caption sdpnal Refer to caption cdcs

### v2

@@ 1 Introduction
    GPU speedup: cuADMM. At each iteration, sGS-ADMM alternates between (a) performing projection onto positive semidefinite (PSD) cones, and (b) solving sparse linear systems. Due to the existence of many small-to-medium-scale PSD cones in the SDP relaxation, we engineer a highly optimized GPU-based implementation of sGS-ADMM in CUDA, named cuADMM and achieves up to 10×10$\times$10 × speedup compared to existing ADMM solvers such as cdcs [78] and sdpnal+ [77]. cuADMM is the first ADMM-based SDP solver that runs in GPUs and we demonstrate its ability to solve five trajectory optimization applications (inverted pendulum, cart-pole, vehicle landing, flying robot, and car back-in with obstacle avoidance): cuADMM solves them with optimality certificates (suboptimality below $1\%$) in minutes (when other solvers take hours and run out of memory) and seconds (when others take minutes).
@@ 2.1 Polynomial Optimization with Chain-like Sparsity
then the following POP is said to admit a chain-like sparsity pattern
@@ 2.2 The Sparse Moment-SOS Hierarchy
Theorem 2.1 provides a way to check the tightness of the relaxation and extract optimal solutions $z^{\star}$ of the nonconvex POP (5). In fact, condition (iii) guarantees the optimal TMS $\varphi^{\star}$ admits a representing Dirac measure supported on the unique optimal solution $z^{\star}$. Therefore, one can extract $z^{\star}$ from the entries of $\varphi^{\star}$ corresponding to degree-one monomials. Note that condition (iii) is a special case of the flat extension condition [46, 56, 16], which states that even when the rank of $M_{k}^{\star}$ is larger than one, say $\mathrm{rank}\left(M_{k}^{\star}\right)=r$, as long as $r$ is consistent with the rank of a submatrix of $M_{k}^{\star}$ and the restriction of moment matrices to each intersection of cliques is rank-one, then global optimality can be certified and exactly $r$ minimizers of the POP can be extracted. We do not detail this setup because in our applications $\mathrm{rank}\left(M_{k}^{\star}\right)$ is always one (up to numerical precision).
@@ 2.3 Faster Conversion to Standard SDP
Solving the moment relaxation (10) is typically done by converting (10) into a standard linear conic optimization problem in the format of sedumi [63] or mosek [6]. The standard way to convert (10) into a conic program follows [74, §5.7.1] (e.g., as in tssos, sostools, yalmip), which will generate a conic program with not only PSD variables but also free (unconstrained) variables.
We advocate for a different conversion standard [74, §5.7.2] that generates only PSD variables and the resulting semidefinite program admits the same chain-like structure as the POP (cf. Fig. 1).
@@
Linear Objective $\left\langle C,X\right\rangle$. Recall the $k$-th clique contains variables $z(I_{k})=(x_{k-1},u_{k-1},x_{k})$. Set $k=2$, with relaxation order $\kappa=2$, the moment matrix (recall it is symmetric and hence we only write the upper triangular part)
@@ 3 Algorithmic Backbone: sGS-ADMM
sGS-ADMM can be seen as a semi-proximal ADMM method applied to the Augmented Lagrangian of (22). Without going deep into the theory, and denoting $\Pi_{\Omega_{+}}(\cdot)$ as the projection onto $\Omega_{+}$, we present sGS-ADMM in Algorithm 1. Global convergence of sGS-ADMM for the SDP pair (13)-(22) is established in [44, 14]. From Algorithm 1, we see two main computational tasks are solving linear systems with a fixed coefficient matrix ${\cal A}{\cal A}^{*}$ and computing the projection of a sequence of symmetric matrices onto the PSD cone. As we shall see in §4, these two tasks can be implemented efficiently in GPUs, resulting in significant speedups.
@@ 5 Experiments
Baselines. For interior point methods, we choose the commercial solver mosek [6] as our baseline. For first-order methods, we choose the state-of-the-art general SDP solvers cdcs [78] and sdpnal+ [77]. [^3] We have tested scs [57] but found its Matlab version to always generate wrong results for multiple-block SDP problems. We only use the ADMM solver in sdpnal+ because the semismooth Newton solver is not scalable. For first-order methods, tol is set as $10^{-4}$ in Algorithm 1.
@@
Summary Results. For the pendulum problem, we tested the performance of four SDP solvers on 100 initial states generated by a dense grid. For the other four experiments, we assessed cuADMM’s performance on 10 randomly generated initial states. Due to the significantly longer solving times of the other three solvers, we evaluated their performance on only one random initial state that cuADMM could solve, with a maximum running time set at 5 hours. Table 1 presents the problem scale, suboptimality gaps, and total running times. It shows that (1) the mean and median suboptimality gaps from cuADMM are either well below $10^{-2}$ (in all cases except for the car back-in) or near $10^{-2}$ (for the car back-in); (2) mosek can solve the pendulum and cart-pole problems very well, but cannot scale to other problems; (3) Compared with other ADMM-based solvers, cuADMM achieves up to 10×10$\times$10 × speedup in large-scale SDP problems (with $m>5\times 10^{5}$); In fact, the other ADMM solvers cannot produce $<1\%$ suboptimality certificates with the time presented in Table 1; (4) With warmstart techniques, cuADMM can solve the pendulum problem to global optimality with an average cost of 0.66 seconds. We only list cuADMM and sdpnal+’s warm start results since cdcs does not support explicit initial guesses.
@@
mosek cdcs sdpnal+ cuADMM
Pendulum $4$ $1$ $55$ $30$ $10$ $47351$ $\bm{-2.58}$ ($\bm{-2.93}$) $9.5$s $110.3$s cold: $264.4$s cold: $31.2$s ($14.6$s)
$\bm{-2.71}$ ($\bm{-2.97}$) warm: $2.95$s warm: $\bm{0.66}$s ($\bm{0.62}$s)
@@
Warmstarts with kNN. To enable data-driven methods for warmstarts, we collect $60\times 120$ mosek solutions in the state space $(\theta,\dot{\theta})\in[0,\pi]\times[-5,5]$ off-line. Given a new initial state $(\theta_{0},\dot{\theta}_{0})$, we use Delaunay Triangulation [41] to search for three nearest neighbors among the mosek’s solutions. A convex combination of these three solutions is treated as the initial solution for cuADMM and sdpnal+.
Comparison with baselines. In the state space $(\theta,\dot{\theta})\in[0,\pi]\times[-5,5]$, we sample $10\times 10$ initial states with a dense grid. Fig. 4 bottom shows the performance of different solvers. In almost all experiments, mosek achieved suboptimality gaps as low as $10^{-7}$. Among the first-order solvers, cdcs struggled to get low dual infeasibility $\eta_{d}$, resulting in significantly higher suboptimality gaps. While sdpnal+ shows reliable performance, its per-iteration cost is $7$ times higher than cuADMM. With warmstarts, cuADMM solves the pendulum problem in less than $0.7$s on average, achieving 10×10$\times$10 × speedup compared to mosek. $10\%$ initial states are hard to solve for all solvers, which form a mysterious spiral line as discussed in [25]. See §0.E.1 for a detailed discussion.
@@
Figure 4: Top: MPC of inverted pendulum with kNN warmstarts. Bottom: performance of six SDP solvers in inverted pendulum.
@@ 0.E.1 Inverted Pendulum
Hyepr-parameters. Denote $x_{k}$ as $\left[r_{c,k}\;r_{s,k}\;f_{s,k}\;f_{s,k}\right]\in{{\mathbb{R}}^{4}}$. In all experiments, $m=1,l=1,b=0.1,g=9.8$ and $N=30,\Delta t=0.1,f_{c,\min}=0.5,u_{\max}=5$. $P_{f}$ in (34) is set to $1$. The number of localizing matrices is $60$. For first-order methods, maxiter is set to $10000$ without kNN warm starts and $100$ with kNN warm starts.
@@
Refer to caption sdpnal+ Refer to caption cdcs